1. What is the concentration of NaCl in a solution if titration of 15.00mL of the solution with 0.2503M AgNO3 requires 20.22mL of the AgNO3 solution to reach the end point?
AgNO3+NaCl->AgCl+NaNO3
2. sometimes leaving the bicycle outside in a hot day will cause tire blowout. why? exolain using KMT
3. what is the vapor pressure of water at 1.00atm and 100’C? using KMT
Answer #1

Here you go:

1) we have to find it out by the formula V1S1=V2S2 where V is the volume and S is the concentration or strength of the solution. Given V1=20.22 mL S1=0.2503M and V2=15 mL. So S2=V1S1/V2 = 20.22×0.2503/15 = 0.3374 M.

According to HOMEWORKLIB RULES I can't answer the rest of the questions and please post it separately and have a good time!!